40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 110     Entry time: Sat Apr 13 21:06:02 2013
Author: Koji 
Type: Optics 
Category: General 
Subject: OMC Bottom-side: cavity glued 

[Jeff, Zach, Lisa, Koji]

Gluing of the cavity mirrors went very well!!!

Preparation

- Checked if the cavity is still resonating. => Yes.

- Checked the FSR: 264.251MHz => 1.1345m
  2.5mm too long => Move each micrometer by 0.625mm backward

- FSR&TMS (I)
  Aligned the cavity again and checked the FSR: 264.8485MHz => 1.13194m
  TMS(V): 58.0875MHz => gamma_V = 0.219324
  TMS(H): 58.1413MHz => gamma_H = 0.219526
  the 9th modes of the carrier is 9.7-10.4 line width (LW) away from the carrier resonance
  the 13th modes of the lower f2 sideband are 9.2-10.2 LW away
  the 19th modes of the upper f2 sideband are 0.3-1.8 LW away
  We found that this coincidence of the resonance can be corrected by shortening the cavity round-trip by 0.5mm

- Spot positions (I)
 
The spots on the curved mirrors were ~1mm too much inside (FM side). In order to translate the cavity axis,
  MM2 and MM4 were pushed by θ
  θ/2.575 = 1mm ==> θ = 2.6 mrad
  The separation of the micrometers are ~20mm
  d/20mm = 2.6mrad ==> d = 52um

  1div of the micrometer corresponds to 10um => 5div = 50um

- Move the micrometers and adjusted the input steering to recover the alignment.

- In any case we were confident to adjust the FSR/TMS/spot positions only with the micrometers

BS1/FM1/FM2 gluing

- Aligned the cavity

- Glued BS1/FM1/FM2 one by one while the cavity resonance was maintained.
  FM2 was slipping as the table is not leveled well and the fixture was not supporting the optic.

- FSR&TMS (II)
  FSR: 264.964875MHz => 1.13144m (Exactly 0.5mm shorter!)
  TMS(V): 58.0225MHz => gamma_V = 0.218982
  TMS(H): 58.1225MHz => gamma_H = 0.219359
  the 9th modes of the carrier is 10.3~11.7 LW away
  the 13th modes of the lower f2 sideband are 7.4~9.3 LW away
  the 19th modes of the upper f2 sideband are 1.5~4.4 LW away

- Spot positions (II)
  Looked OK.

CM2 gluing

- Glued CM2. The mirror was supported from the back with allen keys.

- FSR&TMS (III)
 
FSR: 264.9665625MHz => 1.13144m
  TMS(V): 58.1275MHz => gamma_V = 0.219377
  TMS(H): 58.0813MHz => gamma_H = 0.219202
  the 9th modes of the carrier is 10.2~10.9 LW away
  the 13th modes of the lower f2 sideband are 8.5~9.4 LW away
  the 19th modes of the upper f2 sideband are 1.4~2.7 LW away

- Spot positions (III)
  Looked slightly off at CM2. Pushed MM2 by 4um.

CM1 gluing

- Glued CM1.

- FSR&TMS (IV)
 
FSR: 264.964875MHz => 1.13144m
  TMS(V): 58.06625MHz => gamma_V = 0.219145
  TMS(H): 58.08625MHz => gamma_H = 0.219220
  the 9th modes of the carrier is 10.8~11.1 LW away
  the 13th modes of the lower f2 sideband are 8.2~8.6 LW away
  the 19th modes of the upper f2 sideband are 2.6~3.2 LW away

- Spot positions (final confirmation)
  Looked OK. 

Final measurement

- After everything was finished, more detailed measurement has been done.

- FSR&TMS (final)
 
FSR: 264.963MHz => 1.13145m
  TMS(V): 58.0177MHz => gamma_V = 0.218966
  TMS(H): 58.0857MHz => gamma_H = 0.219221
  the 9th modes of the carrier is 10.8~11.7 LW away
  the 13th modes of the lower f2 sideband are 7.3~8.6 LW away
  the 19th modes of the upper f2 sideband are 2.6~4.5 LW away

Final values for the micrometers

  • MM1: The one closest to the input mirror (CM1) 0.78mm
  • MM2: The other one on CM1 0.89
  • MM3: The one closest to the output mirror (CM2) 0.90
  • MM4: The other one on CM2 0.90
/------------\

0.90         0.78
\------------/
0.90         0.89

 

Attachment 1: Cav_scan_response_130412_Pitch.pdf  203 kB  Uploaded Tue Apr 23 19:35:00 2013  | Hide | Hide all | Show all
Cav_scan_response_130412_Pitch.pdf
Attachment 2: Cav_scan_response_130412_Yaw.pdf  204 kB  Uploaded Tue Apr 23 19:35:11 2013  | Show | Hide all | Show all
ELOG V3.1.3-